OpenStudy (coolkat101):

Can someone please help me on to determine the initial velocity of a marble that is launched horizontally from a CPO marble launcher? it tells me that I must also use formulas that are relevant.. Thank you! ☺

OpenStudy (coolkat101):

@mathstudent55

OpenStudy (coolkat101):

if you could please walk me through how to do this, because i honestly don't know what to do...

OpenStudy (coolkat101):

this is what i wrote so far:

OpenStudy (sapphiremoon):

This seems like projectile motion, so there are three steps, I think. The first one would be to rearrange the initial velocity equation to have the initial velocity as the subject: \[v_{0y} = v_{0}\sin \theta_{0}\] becomes, algebraically, \[v_{0} = \frac{ v_{0y}}{ \sin \theta_{0} }\] So now if we know the firing angle and the initial speed in the y direction we can find the initial velocity. But it's unlikely we'll know the initial y velocity without the initial velocity, so there are ways to find that init. y velocity. How about the second kinematic equation? \[\Delta y = v_{0y}t + \frac{ 1 }{ 2 } g t^{2}\] We can rearrange this to give us v naught y: \[v_{0y} = \frac{ \frac{ 1 }{ 2 }g t^{2} - \Delta y }{ t }\] Now if we have the change in y and time we can get the initial velocity. Does this help? It probably doesn't, but I'm here for questions...

OpenStudy (coolkat101):

I think it helped me! :) And my assignment says to guess the experimental velocity, so would i just assume that the velocity of the hortizontal path of trajetory would be 0 ?

OpenStudy (sapphiremoon):

Not necessarily. The vertical velocity is zero, sorry I didn't realize it was launched horizontally, you'd probably want to change the ys to xes and the sin to cos, the acceleration would be 0 which would eliminate that entire part of the second kinematic equation, and you'd find that: \[v_{0x} = \frac{ \Delta x }{ t}\] Which you knew anyway. :) Horizontal firing totally simplifies the whole thing...

OpenStudy (coolkat101):

oh ok that makes sense! :)

OpenStudy (coolkat101):

sorry for the low quality but this was the original picture of what was assigned if you would like to read off it...

OpenStudy (coolkat101):

well thank you so much and I really appreciate your hard in assisting me with this problem and I'm glad i came to use this website and found someone like you :) Thanks!! I hope have a great day! :)

OpenStudy (coolkat101):

i meant (hard work)

OpenStudy (sapphiremoon):

You're welcome! (Just sayin, if you're new, the best response button is for rewarding the person who helped you the most.)

OpenStudy (coolkat101):

ok there you go! :)